OpenStudy (anonymous):

Four men if working together on a job can complete it in 1 week. If the job is to completed in 4 days, how many more men are required to do the job?

OpenStudy (unklerhaukus):

4m=7d xm=4d

OpenStudy (unklerhaukus):

where x=(n-4)

OpenStudy (unklerhaukus):

solve the first equation for m ie m=4/7 d substitute into the second equation x(4/7 d )=4d x=7 n=3

OpenStudy (anonymous):

thanks (:

OpenStudy (unklerhaukus):

can you decipher all my letters

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a question similar to this, @UnkleRhaukus , Six staff can finish a task in five days. How my days would it take four staff?

OpenStudy (unklerhaukus):

6s=5d 4s=?d

OpenStudy (unklerhaukus):

solve the first equation for s(d)

OpenStudy (unklerhaukus):

i am not going to do it this time but i can tell you what to do if you get stuck

OpenStudy (unklerhaukus):

s=

OpenStudy (anonymous):

errr so that would be s= 6/5?

OpenStudy (anonymous):

and then

OpenStudy (unklerhaukus):

\[s=\frac56 d\]

OpenStudy (unklerhaukus):

substitute s into\[4s=?d\]the divide both sides by d

OpenStudy (unklerhaukus):

\[4\times\frac{5}{6}d=?d\] \[\frac {10}{3}d=?d\] \[\frac {10}{3}=?\]

OpenStudy (anonymous):

ohhh i did that but the multiple choice questions didn't have the answer 3.33

OpenStudy (unklerhaukus):

what are the options?

OpenStudy (anonymous):

6.5 days, 7 days, 8 days or 8.5 days

OpenStudy (anonymous):

i think it should be longer than 5 days....but I'm not sure how to work it out

OpenStudy (unklerhaukus):

you are right it should take more that 5 days

OpenStudy (unklerhaukus):

Six staff can finish a task in five days. \[6s\times5d=t\]How my days would it take four staff?\[4s\times ?d=t\] \[30sd=4?sd\]\[7.5sd=?sd\]\[7.5=?\] still not one of the options i must be doing something wrong

OpenStudy (anonymous):

ahaha, do you know how i can do it or..?

OpenStudy (anonymous):

if 7.5 days is the answer you got, then its good enough for me!

OpenStudy (anonymous):

they probably did it wrong in the book